A company has an EPS of $2.10, a book value per share of $20.58, and a market/book ratio of 3.4x. What is its P/E ratio? The stock price should be rounded to the nearest cent. Round your answer to two decimal places.

x

A firm has a profit margin of 6% and an equity multiplier of 1.9. Its sales are $280 million, and it has total assets of $84 million. What is its ROE? Do not round intermediate calculations. Round your answer to two decimal places.

%
Answer #1
1)
P/E ratio is 33.32
Working:
Step-1:Calculation of market price per share
Market Price per share = Book Value per share * Market/Book ratio
= $    20.58 * 3.4
= $    69.97
Step-2:Calculation of P/E ratio
P/E ratio = Market Price per share/EPS
= $    69.97 / $       2.10
= 33.32
2)
ROE is 10.53%
Working:
Step-1:Calculation of net income
Net Income = Sales * profit margin
= 280*6%
= $    16.80 Million
Step-2:Calculation of stockholder's equity
Stockholder's Equity = Total assets * Equity multiplier
= 84*1.9
= $ 159.60 Million
Step-3:Calculation of return on equity
Return on Equity = Net Income / Stockholder's Equity
= $    16.80 / $ 159.60
= 10.53%